Show that the series summation of 1/a^2 from a=1 to infinity converge and find the sum of the series. Hence or otherwise prove the summation from a=1 to infinity is equal to π^2/6


1
Expert's answer
2021-06-15T18:11:20-0400

The series "\displaystyle\sum_{a=1}^{\infin}\dfrac{1}{a^2}" converges (p-series with "p=2" ).

The Taylor series for the function "f(x)=\sin(x)" is


"\sin(x)=x-\dfrac{x^3}{3!}+\dfrac{x^5}{5!}-\dfrac{x^7}{7!}+..."

Since the equation "f(x)=\sin(x)=0" has the following roots


"x=0, x=\pm\pi, x=\pm2\pi,x=\pm3\pi, ..."

we have


"\sin(x)=Cx(x-\pi)(x+\pi)(x-2\pi)(x+2\pi)..."


"\sin(x)=Cx(x^2-\pi^2)(x^2-2^2\pi^2)(x^2-3^2\pi^2)..."

Then


"\sin(x)=x-\dfrac{x^3}{3!}+\dfrac{x^5}{5!}-\dfrac{x^7}{7!}+..."

"=Cx(x^2-\pi^2)(x^2-2^2\pi^2)(x^2-3^2\pi^2)..."


"1-\dfrac{x^2}{3!}+\dfrac{x^4}{5!}-\dfrac{x^6}{7!}+..."


"=C(x^2-\pi^2)(x^2-2^2\pi^2)(x^2-3^2\pi^2)..."



Substitute "x^2" by "x," we have


"1-\dfrac{x}{3!}+\dfrac{x^2}{5!}-\dfrac{x^3}{7!}+..."

"=C(x-\pi^2)(x-2^2\pi^2)(x-3^2\pi^2)..."

By normalization, we obtain


"1-\dfrac{x}{3!}+\dfrac{x^2}{5!}-\dfrac{x^3}{7!}+..."

"=\dfrac{(x-\pi^2)}{-\pi^2}\cdot\dfrac{(x-2^2\pi^2)}{-2^2\pi^2}\cdot\dfrac{(x-3^2\pi^2)}{-3^2\pi^2}..."

"=(1-\dfrac{x}{\pi^2})(1-\dfrac{x}{2^2\pi^2})(1-\dfrac{x}{3^2\pi^2})..."


Use Vieta formula

Comparing the coefficients of "x"  in both sides of the equation


"-\dfrac{x}{3!}=-\dfrac{x}{\pi^2}-\dfrac{x}{2^2\pi^2}-\dfrac{x}{3^2\pi^2}-..."

Therefore


"-\dfrac{1}{3!}=-\dfrac{1}{\pi^2}-\dfrac{1}{2^2\pi^2}-\dfrac{1}{3^2\pi^2}-..."


"\dfrac{1}{1}+\dfrac{1}{2^2}+\dfrac{1}{3^2}+...=\dfrac{\pi^2}{3!}"

"\displaystyle\sum_{a=1}^{\infin}\dfrac{1}{a^2}=\dfrac{\pi^2}{6}"